OpenStudy (anonymous):

What polynomial has roots of -6, 1, and 4?

OpenStudy (jdoe0001):

\(\large { -6, 1, 4 \begin{cases} x=-6\to x+6=0\to &(x+6)=0\\ x=1\to x-1=0\to &(x-1)=0\\ x=4\to x-4=0\to &(x-4)=0 \end{cases}\qquad thus \\ \quad \\ (x+6)(x-1)(x-4)=\textit{original polynomial} }\)

TheSmartOne (thesmartone):

And if you don't need it factored just multiply it all together and you should be getting something that starts with x^3

OpenStudy (anonymous):

i still don't really understand. here are the answer choices, x3 - 9x2 - 22x + 24 x3 - x2 - 26x - 24 x3 + x2 - 26x + 24 x3 + 9x2 + 14x - 24

TheSmartOne (thesmartone):

It is C...

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

how to i give you a medal? lol

TheSmartOne (thesmartone):

There should be a button next to my response like this http://prntscr.com/4xunzq